A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1- hour flight is .02. What is the probability that (a) both will fail? (b) Neither will fail? (c) one or the other will fail? Show all steps carefully.
: A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1- hour flight is .02. What is the probability that (a) both will fail? (b) Neither will fail? (c) one or the other will fail? Show all steps carefully.

Answer by stanbon(18999)

A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1- hour flight is .02. What is the probability that
(a) both will fail? : (0.02)^2 = 0.0004
------------------------------------------
(b) Neither will fail?: (0.08)^2 = 0.0064
------------------------------------------
(c) one or the other will fail? : 1 - (0.0004 + 0.0064) = 1-0.0068
= 0.9932
==============
Cheers,
Stan H.